Results 1 to 7 of 7

Thread: It used to work just fine!!

  1. #1
    gashead's Avatar
    gashead is offline Admin Verified: Best Accent Ever(Tm)
    Join Date
    May 2009
    Location
    Burntisland, Scotland
    Posts
    887

    Default It used to work just fine!!

    Until I took it apart...

    Greetings,

    Hopefully a quick and easy fix awaits me, here's the story..

    Quite a few years back I bought a 1w 532, (well it was sold as 1w but it actually did about 850, so I knocked the price down)

    This laser and power supply has always run perfectly and when measured was doing around about 800+mw.

    It came with a "Lab style" power supply and because I wanted to rebuild it all into a smaller package I decided to rip it all apart and pack it in a case.

    The months (years) passed and recently I found the time to get it all working again, the problem is.. it doesn't well not to the output power I was expecting.

    When taking the power supply apart I labeled things well, so I knew what went where and so forth, it's all back together as it should apart from the driver is currently without a fan behind it (it had one in the lab case).. seems to get *very* hot.

    Anyway, here is what happens...

    Turn the thing on, power climbs to 600mw, and at that point.. it slows and then drops back from there..

    If I let it cool.. it does the same thing over and over.. 0-600mw.

    600mw "feels" like a ceiling judging by the readings on the meter..

    Tec issue ??

    Questions:

    Does the driver need to be fan cooled ?? this cant be my issue can it ?

    Can I test continuity of the two tec circuits by just putting a multi meter across the + and - ? or should I not do this ?

    The driver itself is TTL which sucks.. should I perhaps look at a flexmod or similar?

    So many questions.

    I'll provide pics and perhaps youtube.. but not now.. I'm sad and tired.. work sucked today and now my beautiful Laser is sick.

    Any help / advice / condolences welcome.

    Cheers,

    Nige.

  2. #2
    Join Date
    Feb 2010
    Location
    Ca.
    Posts
    84

    Default

    Ok,I'll probably be flogged and chastized for suggesting this,but it has worked for me in the past.
    (no joke by the way)
    Put the head in the freezer for about 5 minutes,remove and power up whilst noting output.
    Should be low or none until the IR heats up the magic crystals enough to start lasing.
    If it does,note the temp. of the holder at peak output and experiment from there with TEC setting,no?
    Worked for me trying to determine mode hopping and output variations on DPSS pointers.
    And yes your ps should have fan working,unless you want to see other problems.
    My opinion of course!

  3. #3
    Join Date
    Jul 2005
    Location
    Flying over a town near you
    Posts
    1,404

    Default

    Temperature issues it would seem...any pix of the head?
    You are the only one that can make your dreams come true....and the only one that can stop them...A.M. Dietrich

  4. #4
    gashead's Avatar
    gashead is offline Admin Verified: Best Accent Ever(Tm)
    Join Date
    May 2009
    Location
    Burntisland, Scotland
    Posts
    887

    Unhappy

    Quote Originally Posted by MechEng3 View Post
    Temperature issues it would seem...any pix of the head?


    Click image for larger version. 

Name:	2010-09-21 06.32.37.jpg 
Views:	16 
Size:	390.0 KB 
ID:	19716Attachment 19717

    Sorry about the poor lighting conditions - shot at 06:30 on my camera phone, in my workshop where my Wife thought it would be a good idea to remove the 100watt bulb and replace with a swankey Ikea lampshade (Slightly yellow) and 40 watt bulb.

    I'm more and more convinced that I have a break in one of the wires leading to one of the tec's - this would explain the behavior I guess..

    Cheers!

    N.

  5. #5
    gashead's Avatar
    gashead is offline Admin Verified: Best Accent Ever(Tm)
    Join Date
    May 2009
    Location
    Burntisland, Scotland
    Posts
    887

    Default

    Quote Originally Posted by HD58PHD View Post
    Ok,I'll probably be flogged and chastized for suggesting this,but it has worked for me in the past.
    (no joke by the way)
    Put the head in the freezer for about 5 minutes,remove and power up whilst noting output.
    Should be low or none until the IR heats up the magic crystals enough to start lasing.
    If it does,note the temp. of the holder at peak output and experiment from there with TEC setting,no?
    Worked for me trying to determine mode hopping and output variations on DPSS pointers.
    And yes your ps should have fan working,unless you want to see other problems.
    My opinion of course!

    I won't flog / nay chastize you! Help and tips are what makes this community great!

    I'm avoiding touching the TEC setting, as the driver is tuned for the head, everything worked before I pulled it apart, so I'm guessing it's a electrical issue. (Ie: something is broken)

    As you can see from pics, Power supply (and driver) is not restricted with airflow - plan is to box this all up and add fans for movement of air - but I'd be surprised if no fan = this problem... could be wrong though!

    N.

  6. #6
    Join Date
    Jul 2005
    Location
    Flying over a town near you
    Posts
    1,404

    Default

    Retrace your thermistor signal wires and be certain of a correct reading on them...the resisitance. Check the IRL150 that drives the TECs also. I have seen a lot of strange things go wrong. Unless you go inside the head to set the temps with a pair of digtal temp probes.....I would leave it....but if you could pop the head open to see if it was near the correct temps, it may help to lock in the problems. Beware of multi watt IR if you do.
    You are the only one that can make your dreams come true....and the only one that can stop them...A.M. Dietrich

  7. #7
    gashead's Avatar
    gashead is offline Admin Verified: Best Accent Ever(Tm)
    Join Date
    May 2009
    Location
    Burntisland, Scotland
    Posts
    887

    Default

    Quote Originally Posted by MechEng3 View Post
    Retrace your thermistor signal wires and be certain of a correct reading on them...the resisitance. Check the IRL150 that drives the TECs also. I have seen a lot of strange things go wrong. Unless you go inside the head to set the temps with a pair of digtal temp probes.....I would leave it....but if you could pop the head open to see if it was near the correct temps, it may help to lock in the problems. Beware of multi watt IR if you do.

    Fixed! - Embarrassing user error - too many late nights mixed with two children under two years old.

    (Four wires from the D Sub connector on the driver were not connected, assuming they are the thermistor signals.. felt really stupid, reconnected, get stable 758mW after optics... happy again!)

    Cheers for the help MechEng, owe you beer!



    N.

Posting Permissions

  • You may not post new threads
  • You may not post replies
  • You may not post attachments
  • You may not edit your posts
  •